OpenStudy (anonymous):

Solve for x: x - 1 ----- = 2 9

OpenStudy (nubeer):

quite dont understand the format of your question.. can u retype it and this time in better format?

hartnn (hartnn):

multiply 9 on both sides, what u get ?

hartnn (hartnn):

\(\large \frac{x-1}{9}=2\)

hartnn (hartnn):

\(\large \frac{x-1}{9}\times 9=2 \times 9\) can u tell me next step ?

hartnn (hartnn):

NO?

Parth (parthkohli):

lol